### v1

# On the Localization of the Christoffel-Darboux Kernel

Abstract. We establish that the Christoffel-Darboux (or Bergman) kernel localizes along the diagonal for non-pluripolar measures. This leads to several applications in the theory of Toeplitz operators, particularly concerning their asymptotic algebra property and spectral equidistribution.
1 Introduction
The primary goal of this paper is to establish a diagonal localization phenomenon for the Christoffel-Darboux kernel associated with a broad class of measures.
To state our results precisely, we fix a compact complex manifold $X$, $\dim_{\mathbb{C}}X=n$, and a big line bundle $L$ over $X$, i.e. such that for some $c>0$, we have $\dim H^{0}(X,L^{\otimes k})\geq ck^{n}$, for $k\in\mathbb{N}$ big enough. We fix a continuous Hermitian metric $h^{L}$ on $L$ and a non-pluripolar Borel measure $\mu$ on $X$, i.e. such that for any pluripolar subset $E\subset X$, we have $\mu(E)=0$. Recall that a subset $E\subset X$ is called pluripolar if in each local chart we have $E\subset\{\phi=-\infty\}$ for some plurisubharmonic (psh) function $\phi$. We denote by ${\textrm{Hilb}}_{k}(h^{L},\mu)$ the positive semi-definite form on $H^{0}(X,L^{\otimes k})$ defined for arbitrary $s_{1},s_{2}\in H^{0}(X,L^{\otimes k})$ as follows
$\langle s_{1},s_{2}\rangle_{{\textrm{Hilb}}_{k}(h^{L},\mu)}=\int\langle s_{1}(x),s_{2}(x)\rangle_{(h^{L})^{k}}\cdot d\mu(x).$ (1.1)
Remark that since $\mu$ is non-pluripolar, the above form is positive definite.
We assume that $k$ is large enough so that $n_{k}>0$, where $n_{k}\mathrel{\mathop{\ordinarycolon}}=\dim H^{0}(X,L^{\otimes k})$. The central object of this article is the Christoffel-Darboux (or Bergman) kernel, [1], defined as
$B_{k}(x,y)\mathrel{\mathop{\ordinarycolon}}=\sum_{i=1}^{n_{k}}s_{i}(x)\cdot s_{i}(y)^{*}\in L^{k}_{x}\otimes(L^{k}_{y})^{*},$ (1.2)
where $s_{i}$, $i=1,\ldots,n_{k}$, is an orthonormal basis of $(H^{0}(X,L^{\otimes k}),{\textrm{Hilb}}_{k}(h^{L},\mu))$ (an easy verification shows that (1.2) doesn’t depend on the choice of the basis).
To grasp the intuitive meaning behind the Christoffel-Darboux kernel, we denote by $\Sigma_{k}$ the base loci of $L^{k}$, i.e. $\Sigma_{k}\mathrel{\mathop{\ordinarycolon}}=\{x\in X\mathrel{\mathop{\ordinarycolon}}s(x)=0\text{ for any }s\in H^{0}(X,L^{\otimes k})\}$. For $x\in X\setminus\Sigma_{k}$, we denote by $s_{k,x}\in H^{0}(X,L^{\otimes k})$ the peak section at $x$ with respect to the scalar product ${\textrm{Hilb}}_{k}(h^{L},\mu)$; recall that this means that $s_{k,x}$ is the unique section (up to a unimodular complex multiplication) which maximizes the ratio $|s(x)|_{(h^{L})^{k}}/\|s\|_{{\textrm{Hilb}}_{k}(h^{L},\mu)}$ among all $s\in H^{0}(X,L^{\otimes k})\setminus\{0\}$.
Directly from the definition of the Christoffel-Darboux kernel and the fact that it doesn’t depend on the choice of an orthonormal basis, we deduce
$B_{k}(x,y)=\begin{cases}s_{k,x}(x)\cdot s_{k,x}(y)^{*},\quad&x\in X\setminus\Sigma_{k},y\in X,\\ 0,&\text{otherwise.}\end{cases}$ (1.3)
The description of peak sections suggests that, as $k\to\infty$, the values $y\mapsto|s_{k,x}(y)|_{(h^{L})^{k}}$ become increasingly concentrated near $y=x$ for any fixed $x\in X$.
For certain “regular” measures, this heuristic holds in a strong form; see Remark 1.2 for a brief survey. In general, however, naive versions of the phenomenon break down: one does not have pointwise convergence to zero [14, §2], nor even weak convergence of the norm at individual points in the support of $\mu$ [9]; see also Section 6. The main result of this article shows, nevertheless, that the heuristic can be recovered in an averaged form, expressed via the Christoffel-Darboux kernel.
Before turning to the general formulation of our results, we illustrate a connection with orthogonal polynomials. Let $\mu$ be a Borel measure on $\mathbb{C}$ so that $K\mathrel{\mathop{\ordinarycolon}}={\textrm{supp}}(\mu)$ is non-finite and compact. Consider a sequence $p_{i}(z)$, $i\in\mathbb{N}$, $z\in\mathbb{C}$, of orthogonal polynomials associated with $\mu$, cf. [29]. Recall that it means that each $p_{i}$ is a polynomial of degree $i$ in $z$, normalized to have unit norm in $L^{2}(\mu)$, and orthogonal to all $p_{j}$ with $j<i$. The non-finiteness of the support of $\mu$ assures that the scalar product (1.1) is positive definite and hence $p_{j}$ are well-defined (up to a multiplication by a unimodular constant). The associated Christoffel-Darboux kernel, [11], [13], is defined by
$B_{k}^{{\rm{ort}}}(x,y)\mathrel{\mathop{\ordinarycolon}}=\sum_{i=0}^{k-1}p_{i}(x)\cdot\overline{p_{i}(y)},\qquad x,y\in\mathbb{C}.$ (1.4)
To realize this Christoffel-Darboux kernel as a special case of the preceding construction, we consider the projective space $X\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}^{1}$ and the hyperplane bundle $L=\mathscr{O}(1)$. We view $\mathbb{C}$ as an affine chart in $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$ with the standard holomorphic coordinate $z$, and consider a non-zero section $\sigma\in H^{0}(X,L)$ vanishing at $\infty$. The division by $\sigma^{k}$ gives an isomorphism between $H^{0}(X,L^{\otimes k})$ and the space of polynomials of degree $\leq k-1$ in $z$. If we now endow $L$ with a continuous metric $h^{L}$, such that $|\sigma(x)|_{h^{L}}=1$ for all $x\in K$, then the $L^{2}$-norm (1.1) coincides, under the above identification, with the $L^{2}(\mu)$-norm on the space of polynomials. In particular, (1.2) then specializes to (1.4).
We now come back to our general setting of a complex manifold and a big line bundle. We introduce the following sequence of measures on $X\times X$:
$\mu_{k}^{\rm{Berg}}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{n_{k}}|B_{k}(x,y)|^{2}_{(h^{L})^{k}}\cdot d\mu(x)d\mu(y),\quad k\in\mathbb{N}.$ (1.5)
It follows immediately from (1.2) that each $\mu_{k}^{\rm{Berg}}$ is a probability measure. The following result confirms the heuristics above, and it is the main result of this paper.
Theorem 1.1.
For any non-pluripolar Borel measure $\mu$, the corresponding measures $\mu_{k}^{\rm{Berg}}$ do not asymptotically place mass away from the diagonal. In other words, for any compact subset $K\subset X\times X$, not intersecting the diagonal, we have
$\lim_{k\to\infty}\iint_{K}\mu_{k}^{\rm{Berg}}=0.$ (1.6)
Remark 1.2.
There is a large body of work which establishes versions of Theorem 1.1 for some specific measures. A comprehensive survey would take us too far afield, so we merely cite a few representative results. In the complex geometric setting – where the measure arises from a volume form – see the works of Dai-Liu-Ma [12], Christ [10], Ma-Marinescu [25], Berman [2]. For measures supported on totally real algebraic submanifolds, see Berman-Ortega-Cerdá [4]. In the context of orthogonal polynomials, we refer to Nevai [26], Lubinsky [22], Breuer-Last-Simon [9].
We note that while Theorem 1.1 appears to be new even in the context of orthogonal polynomials, the proof of it simplifies considerably in this special case (only Section 5 is relevant), and in fact a stronger statement holds under weaker assumptions: according to Remark 5.2, for arbitrary compactly supported measures $\mu$ on $\mathbb{C}$ with non-finite support and a compact $K\subset\mathbb{C}\times\mathbb{C}$, not intersecting the diagonal, there is $C>0$, so that in the notations of (1.4), for any $k\in\mathbb{N}$, we have
$\iint_{K}\big{|}B_{k}^{{\rm{ort}}}(x,y)\big{|}^{2}\cdot d\mu(x)d\mu(y)\leq C.$ (1.7)
For Bernstein-Markov measures (see [7] for a survey on the Bernstein-Markov coondition), Theorem 1.1 was recently established by the author in [14], building on earlier work of Berman-Boucksom-Witt Nyström [3], cf. also Bloom-Levenberg [5], [6], concerning the equidistribution of the diagonal Christoffel-Darboux measures (1.9), and thereby resolving a question left open by Zelditch [30] concerning the weak convergence of $\mu_{k}^{\rm{Berg}}$. The main objective of this article is to demonstrate that, by employing alternative techniques, it is possible to bypass the equidistribution result [3] and thus work within strictly weaker assumptions.
Our primary motivation for Theorem 1.1 lies in the theory of Toeplitz operators. Let us recall the necessary notations: for $f\in\mathscr{C}^{0}(X)$, $k\in\mathbb{N}^{*}$, we define the Toeplitz operator $T_{k}(f)\in{\text{End}{(H^{0}(X,L^{\otimes k}))}}$ as $T_{k}(f)\mathrel{\mathop{\ordinarycolon}}=B_{k}\circ M_{k}(f)$, where $B_{k}\mathrel{\mathop{\ordinarycolon}}\mathscr{C}^{0}(X,L^{\otimes k})\to H^{0}(X,L^{\otimes k})$ is the orthogonal (Bergman) projection to $H^{0}(X,L^{\otimes k})$ with respect to (1.1), and $M_{k}(f)\mathrel{\mathop{\ordinarycolon}}H^{0}(X,L^{\otimes k})\to\mathscr{C}^{0}(X,L^{\otimes k})$ is the multiplication map by $f$.
One of the key features of Toeplitz operators is their asymptotic algebra property: for $f,g\in\mathscr{C}^{0}(X)$, the composition $T_{k}(f)\circ T_{k}(g)$ is asymptotically close (with respect to some norm) to $T_{k}(f\cdot g)$, as $k\to\infty$. This phenomenon appears to have been first observed by Grenander-Szegő in [17, §7.4] in the context of certain orthogonal polynomials. It was later established in the complex-geometric setting by Bordemann-Meinrenken-Schlichenmaier [8]; see also Ma-Marinescu [24]. In the more general setting of measures verifying the Bernstein-Markov assumption, the result was recently established by the author in [14, §3]. As we shall see below, and somewhat unexpectedly, even the Bernstein-Markov assumption can in fact be dispensed with.
To describe this precisely, it is known that a version of the above asymptotic algebra property for Toeplitz operators is a consequence of the localization of the Christoffel-Darboux kernel, cf. [14, §3]. Repeating the proof of [14, Theorem 1.5], with [14, Theorem 2.1] replaced by the more general statement, Theorem 1.1, immediately yields the following result.
Corollary 1.3.
For any non-pluripolar measure $\mu$, the space of Toeplitz operators verifies the asymptotic algebra property. In other words, for any $f,g\in\mathscr{C}^{0}(X)$, $p\in[1,+\infty[$, we have
$\lim_{k\to\infty}\|T_{k}(f)\circ T_{k}(g)-T_{k}(f\cdot g)\|_{p}=0,$ (1.8)
where $\|\cdot\|_{p}$ is the $p$-Schatten norm, defined for an operator $A\in{\text{End}{(V)}}$, of a finitely-dimensional Hermitian vector space $(V,H)$ as $\|A\|_{p}=(\frac{1}{\dim V}{\rm{Tr}}[|A|^{p}])^{\frac{1}{p}}$, $|A|\mathrel{\mathop{\ordinarycolon}}=(AA^{*})^{\frac{1}{2}}$.
Remark 1.4.
a) It would be interesting to classify those measures for which the $p$-Schatten norm above could be replaced by the operator norm.
b) Combining the above with (1.7), we find that in the setting of orthogonal polynomials, (1.8) holds under weaker assumptions: for arbitrary measures with non-finite support.
As an application of Corollary 1.3, we show in Section 7 that the spectral measures of Toeplitz operators equidistribute if and only if the associated diagonal Christoffel-Darboux measures (1.9) do. This result refines several classical versions of Szegő’s first limit theorem [28], as Boutet de Monvel-Guillemin [20]; see also Ma-Marinescu [23].
To describe the organization of this article, we briefly survey the main steps of the proof of Theorem 1.1. As in the earlier work [14], the key idea – albeit perhaps counterintuitive – is to investigate the off-diagonal behavior of Christoffel-Darboux kernel by analyzing its diagonal behaviour. For this, we define the diagonal Christoffel-Darboux measures on $X$ as
$\mu_{k}^{\rm{Berg},\Delta}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{n_{k}}|B_{k}(x,x)|\cdot d\mu(x).$ (1.9)
Beyond this, our approach departs significantly from [14]: in the absence of the Bernstein-Markov property, little can be said about the equidistribution of (1.9), a feature that was crucial in loc. cit.
To remedy this, we establish in Sections 2 and 3, that any subsequential limit of the sequence of measures $\mu_{k}^{\rm{Berg},\Delta}$, $k\in\mathbb{N}$, assigns zero mass to subvarieties. Using this, by relying on the estimates of Lubinsky [22], we then show in Section 4 that it suffices to establish Theorem 1.1 for measures with support outside of a fixed subvariety. The proof will then be completed in Section 5, where we establish that for measures supported away from some divisors on $X$, a more precise version of Theorem 1.1 holds.
In Section 6, we discuss a pointwise version of our localization result in connection with [9, Conjecture 1.4]. In Section 7, we discuss the applications to the theory of Toeplitz operators.
Acknowledgement. The author thanks Norm Levenberg for several related discussions and Robert Berman for drawing attention to [4]. This work was supported by the CNRS, École Polytechnique, and in part by the ANR projects QCM (ANR-23-CE40-0021-01), AdAnAr (ANR-24-CE40-6184) and STENTOR (ANR-24-CE40-5905-01).
2 Christoffel-Darboux measures put no mass on subvarieties
We retain the notation from Introduction and assume, from now on and throughout the rest of the article, that the measure $\mu$ is non-pluripolar.
The primary objective of this section is to demonstrate that every subsequential limit of the diagonal Christoffel-Darboux measures vanishes on subvarieties.
Theorem 2.1.
For any subvariety $Y\subset X$ and any $\epsilon>0$, there exists an open neighborhood $U$ of $Y$, so that for any $k\in\mathbb{N}$, we have
$\int_{U}\mu_{k}^{\rm{Berg},\Delta}\leq\epsilon.$ (2.1)
Remark 2.2.
Note that equilibrium measure puts no mass on pluripolar subsets, when $\mu$ is Bernstein-Markov. Hence, if $\mu$ is Bernstein-Markov, Theorem 2.1 follows from the equidistribution result of [3] showing the weak convergence of $\mu_{k}^{\rm{Berg},\Delta}$ towards the equilibrium measure.
In order to prove Theorem 2.1, we need to consider the so-called partial Christoffel-Darboux kernels associated with $Y$, denoted by $B_{k}^{\epsilon,Y}(x,y)\in L^{k}_{x}\otimes(L^{k}_{y})^{*}$ for $\epsilon>0$, and defined as
$B_{k}^{\epsilon,Y}(x,y)\mathrel{\mathop{\ordinarycolon}}=\sum_{i=1}^{n_{k}(\epsilon)}s_{i}(x)\cdot s_{i}(y)^{*},$ (2.2)
where $s_{i}$, $i=1,\ldots,n_{k}(\epsilon)$ is an orthonormal basis of $(H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon k\rceil}),{\textrm{Hilb}}_{k}(h^{L},\mu))$, $n_{k}(\epsilon)\mathrel{\mathop{\ordinarycolon}}=\dim H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon k\rceil})$, and $\mathcal{J}_{Y}$ is the sheaf of holomorphic sections vanishing along $Y$. The following result, out of independent interest, will be established in Section 3.
Theorem 2.3.
For any $\epsilon,\delta>0$, there is an open neighborhood $U$ of $Y$, so that for any $x\in U$, $k\in\mathbb{N}$, we have $|B_{k}^{\epsilon,Y}(x,x)|\leq\exp(-\delta k)$.
Remark 2.4.
When $\mu$ is a volume form, Theorem 2.3 follows from more precise estimates of Berman [2, Lemma 4.1].
Proof of Theorem 2.1.
We denote by ${\rm vol}(L)$ the volume of a line bundle $L$ on $X$, defined by
${\rm{vol}}(L)\mathrel{\mathop{\ordinarycolon}}=\limsup_{k\to\infty}\frac{n!}{k^{n}}\dim H^{0}(X,L^{\otimes k}).$ (2.3)
By Fujita’s theorem [16], the above upper limit is actually a limit. The definition of the volume functional naturally extends to the $\mathbb{Q}$-line bundles, and the resulting functional extends continuously to the full Néron-Severi space $NS^{1}(X)$, cf. [21, Theorem 2.2.44].
Remark also that for any $l\in\mathbb{N}$, there is an isomorphism
$H^{0}(\hat{X},\pi^{*}L^{\otimes k}\otimes\mathscr{O}(-l\cdot E))\to H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{l}),$ (2.4)
where $\pi\mathrel{\mathop{\ordinarycolon}}\hat{X}\to X$ is the blow-up along the ideal sheaf $\mathcal{J}_{Y}$ of holomorphic functions vanishing along $Y$, and $\mathscr{O}_{\hat{X}}(-E)\mathrel{\mathop{\ordinarycolon}}=\pi^{*}\mathcal{J}_{Y}$. In particular, the limit $\lim_{k\to\infty}\frac{n!}{k^{n}}\dim H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon^{\prime}k\rceil})$ exists and can be interpreted as the volume (on $\hat{X}$) or the $\mathbb{R}$-line bundle $L\otimes\mathscr{O}_{\hat{X}}(-E)^{\epsilon^{\prime}}$.
From this and the continuity of the volume, for a given $\epsilon>0$, we can find $\epsilon^{\prime}>0$, so that
$\liminf_{k\to\infty}\frac{\dim H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon^{\prime}k\rceil})}{\dim H^{0}(X,L^{\otimes k})}\geq 1-\frac{\epsilon}{2}.$ (2.5)
For the choice $\epsilon\mathrel{\mathop{\ordinarycolon}}=\epsilon^{\prime}$ and $\delta\mathrel{\mathop{\ordinarycolon}}=1$, we now consider a neighborhood $U$ of $Y$ given by Theorem 2.3, so that for any $x\in U$, $k\in\mathbb{N}$, we have
$|B_{k}^{\epsilon^{\prime},Y}(x,x)|\leq\exp(-k).$ (2.6)
By writing $B_{k}(x,x)=\big{(}B_{k}(x,x)-B_{k}^{\epsilon^{\prime},Y}(x,x)\big{)}+B_{k}^{\epsilon^{\prime},Y}(x,x)$, and using the elementary identities
$\int B_{k}(x,x)\cdot d\mu(x)=\dim H^{0}(X,L^{\otimes k}),$ (2.7)
$\int B_{k}^{\epsilon^{\prime},Y}(x,x)\cdot d\mu(x)=\dim H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon^{\prime}k\rceil}),$
together with (2.5), we deduce the bound (2.1) for such $U$ directly from (2.6). ∎
3 Forbidden region for the partial Christoffel-Darboux kernels
The main goal of this section is to establish that the partial Christoffel-Darboux kernels decrease exponentially in the neighborhood of the subvariety, that is, to prove Theorem 2.3. The proof relies on the use of several psh envelopes constructions.
More specifically, we reduce first the estimate of the Christoffel-Darboux kernel associated with the $L^{2}(\mu)$-norm to a related estimate for the $L^{\infty}$-norm. For this, we rely on the non-negligible psh envelopes introduced in [18] and further studied in [15]. Then we apply the tautological maximum principle to relate this estimate with psh envelopes associated with divisors introduced in [2]. Theorem 2.3 then follows from some estimates on such envelopes.
To simplify the presentation, we claim that it suffices to establish Theorem 2.3 under the assumption that $\mathcal{J}_{Y}=\mathscr{O}_{\hat{X}}(-E)$ for some effective divisor $E$ on $X$. To see this, let $\pi\mathrel{\mathop{\ordinarycolon}}\hat{X}\to X$ denote the blow-up of $X$ along $\mathcal{J}_{Y}$, and let $\mathscr{O}_{\hat{X}}(-E)\mathrel{\mathop{\ordinarycolon}}=\pi^{*}\mathcal{J}_{Y}$. If $\hat{X}$ is not smooth, we perform further blow-ups to resolve its singularities. By an abuse of notation, we denote the resolution by $\hat{X}$. Let $\hat{\mu}$ be a measure on $\hat{X}$ such that $\pi_{*}\hat{\mu}=\mu$. Since $\mu$ is non-pluripolar and $\pi$ is a biholomorphism outside an analytic (hence pluripolar) subset, the measure $\hat{\mu}$ is uniquely defined and remains non-pluripolar. The isomorphism (2.4) is then an isometry with respect to the induced $L^{2}$-norms. In particular, the associated diagonal Christoffel-Darboux measures $\hat{\mu}_{k}^{\mathrm{Berg},\Delta}$ on $\hat{X}$ and $\mu_{k}^{\mathrm{Berg},\Delta}$ on $X$ satisfy the compatibility relation $\pi_{*}\hat{\mu}_{k}^{\mathrm{Berg},\Delta}=\mu_{k}^{\mathrm{Berg},\Delta}$. It follows that it is enough to prove Theorem 2.3 on $\hat{X}$. From now on, we assume $\mathcal{J}_{Y}=\mathscr{O}_{\hat{X}}(-E)$ for some effective divisor $E$ on $X$.
Our proof of Theorem 2.3 will be based on the following three auxiliary statements. For the first one, we use the following notation: for a continuous metric $h^{L}_{0}$ on $L$, we denote by ${\textrm{Ban}}_{k}^{\infty}(h^{L}_{0})$ the $L^{\infty}$-norm on $H^{0}(X,L^{\otimes k})$ associated with $h^{L}_{0}$. We also denote $\omega\mathrel{\mathop{\ordinarycolon}}=2\pi c_{1}(L,h^{L})$.
Lemma 3.1.
There is a continuous metric $h^{L}_{0}$ on $L$, so that for any $k\in\mathbb{N}$, we have
${\textrm{Hilb}}_{k}(h^{L},\mu)\geq{\textrm{Ban}}_{k}^{\infty}(h^{L}_{0}).$ (3.1)
Proof.
Following [18], we define the non-negligible psh envelope as
$\phi_{\mu}\mathrel{\mathop{\ordinarycolon}}=\sup\Big{\{}\phi_{0}\in{\textrm{PSH}}(X,\omega)\mathrel{\mathop{\ordinarycolon}}\phi_{0}\leq 0\text{ $\mu$-almost everywhere}\Big{\}},$ (3.2)
where ${\textrm{PSH}}(X,\omega)$ stands for the space of $\omega$-psh functions $\phi\mathrel{\mathop{\ordinarycolon}}X\to[-\infty,+\infty[$; those are upper semi-continuous functions so that $\omega+\sqrt{-1}\partial\overline{\partial}\phi$ is positive as a $(1,1)$-current. It was established by Guedj-Lu-Zeriahi [18, Proposition 2.10], cf. also [15, Proposition 3.1], that $\phi_{\mu}$ is bounded from above on $X$. To make this article self-contained, we recall briefly the argument from [15]. Following Siciak [27] and Guedj-Zeriahi [19], for a subset $E$, we define
$\phi_{E}=\sup\Big{\{}\phi_{0}\in{\textrm{PSH}}(X,\omega)\mathrel{\mathop{\ordinarycolon}}\phi_{0}\leq 0\text{ over }E\Big{\}}.$ (3.3)
It is classical that if $E$ is non-pluripolar, then $\phi_{E}$ is bounded from above, cf. [19, Theorem 9.17].
Immediately from our definitions, we have $\phi_{\mu}=\sup\{\phi_{X\setminus F}\mathrel{\mathop{\ordinarycolon}}\mu(F)=0\}$. Suppose now, for contradiction, that $\phi_{\mu}$ is not bounded from above on $X$. Then there exists a sequence of sets $E_{i}$, $i\in\mathbb{N}$, with $\mu(E_{i})=0$, such that $\sup_{X}\phi_{X\setminus E_{i}}\to+\infty$, as $i\to\infty$. Then for $E\mathrel{\mathop{\ordinarycolon}}=\cup_{i\in\mathbb{N}}E_{i}$, we obviously have $\phi_{X\setminus E_{i}}\leq\phi_{X\setminus E}$ for any $i\in\mathbb{N}$, and as a consequence, $\sup_{X}\phi_{X\setminus E}=+\infty$. However, since $\mu(E)=0$, we have $\mu(X\setminus E)=1$, and so $X\setminus E$ is non-pluripolar, as we have assumed that $\mu$ is non-pluripolar. But then from [19, Theorem 9.17], $\sup_{X}\phi_{X\setminus E}<+\infty$, giving us the required contradiction.
From this point, the proof of Lemma 3.1 follows immediately from the following statement of [15, Theorem 7.1]: for any continuous function $\phi_{0}\mathrel{\mathop{\ordinarycolon}}X\to\real$, verifying $\phi_{\mu}\leq\phi_{0}$, the following holds. For any $\epsilon>0$, there is $C>0$, such that for any $k\in\mathbb{N}$, we have
${\textrm{Hilb}}_{k}(h^{L},\mu)\geq C\cdot\exp(-\epsilon k)\cdot{\textrm{Ban}}_{k}^{\infty}(h^{L}\cdot\exp(-\phi_{0})).$ (3.4)
Note that [15, Theorem 7.1] was stated for ample line bundles. However, since the analytic arguments essentially rely on [3], which applies to big line bundles, the conclusions remain valid in this more general setting. ∎
We now decompose the divisor $E$ into irreducible components $E=\sum_{i=1}^{r}a_{i}E_{i}$, $a_{i}\in\mathbb{N}$. The next lemma is a version of a well-known tautological maximal principle for the $L^{\infty}$-norm. The only minor difference is that we apply it for sections which are vanishing along $E$.
For this, recall that for a psh function $\phi$, defined in a neighborhood of $a\in\mathbb{C}^{n}$, the Lelong number $\nu_{a}(\phi)$ of $\phi$ at $a$ is defined as
[…]
For an irreducible subvariety $Z\subset\mathbb{C}^{n}$, we define the generic Lelong number of $\phi$ along $Z$ as
$\nu_{Z}(\phi)=\inf_{x\in Z}\nu_{x}(\phi).$ (3.6)
We now define the following envelope
$\phi^{\epsilon,E}\mathrel{\mathop{\ordinarycolon}}=\sup\Big{\{}\phi\in{\textrm{PSH}}(X,\omega)\mathrel{\mathop{\ordinarycolon}}\phi\leq 0\text{, and }\nu_{E_{i}}(\phi)\geq\epsilon a_{i},\text{ for any $i=1,\ldots,r$}\Big{\}}.$ (3.7)
Lemma 3.2.
The norms ${\textrm{Ban}}_{k}^{\infty}(h^{L})$ and ${\textrm{Ban}}_{k}^{\infty}(h^{L}\cdot\exp(-\phi^{\epsilon,E}))$ coincide when restricted to the subspace $H^{0}(X,L^{\otimes k}\otimes\mathscr{O}_{X}(-\lceil\epsilon kE\rceil))\subset H^{0}(X,L^{\otimes k})$.
Proof.
Note first that since $\phi^{\epsilon,E}\leq 0$, we immediately get the inequality ${\textrm{Ban}}_{k}^{\infty}(h^{L})\leq{\textrm{Ban}}_{k}^{\infty}(h^{L}\cdot\exp(-\phi^{\epsilon,E}))$. To prove Lemma 3.2, it remains to establish the reverse inequality.
For a given section $s\in H^{0}(X,L^{\otimes k}\otimes\mathscr{O}_{X}(-\lceil\epsilon kE\rceil))\setminus\{0\}$, we consider the function $\phi(x)\mathrel{\mathop{\ordinarycolon}}=\frac{1}{k}\log|s(x)|_{(h^{L})^{k}}-\frac{1}{k}\log\|s\|_{{\textrm{Ban}}_{k}^{\infty}(h^{L})}$. By the Poincaré-Lelong formula, $\phi$ is one of the admissible candidates in the supremum defining (3.7). The obvious inequality $\phi(x)\leq\phi^{\epsilon,E}(x)$ is simply a restatement of the pointwise bound $|s(x)|_{(h^{L}\cdot\exp(-\phi^{\epsilon,E}))^{k}}\leq\|s\|_{{\textrm{Ban}}_{k}^{\infty}(h^{L})}$. Taking the supremum over $x\in X$ in the above inequality yields the desired conclusion. ∎
For the last auxiliary result, we denote by $s_{E_{i}}\in H^{0}(X,\mathscr{O}_{X}(E_{i}))$ the canonical holomorphic section verifying ${\textrm{div}}(s_{E_{i}})=E_{i}$. We endow each line bundle $\mathscr{O}_{X}(E_{i})$, $i=1,\ldots,r$, with a smooth metric $h^{E_{i}}$, and denote by $x\mapsto|s_{E_{i}}(x)|$ the norms of the associated sections.
Lemma 3.3.
Assume that $\epsilon>0$ is small enough so that $c_{1}(L)-\epsilon\cdot\sum_{i=1}^{r}a_{i}\cdot c_{1}(\mathscr{O}_{X}(E_{i}))$ is a big class (i.e. containing a strictly positive $(1,1)$-current). Then there is $C>0$ so that for any $x\in X$,
$\phi^{\epsilon,E}(x)\leq\epsilon\cdot\sum_{i=1}^{r}a_{i}\cdot\log|s_{E_{i}}(x)|+C.$ (3.8)
Proof.
Recall that Berman in [2, Lemma 4.1] established the following identity
$\phi^{\epsilon,E}-\epsilon\cdot\sum_{i=1}^{r}a_{i}\cdot\log|s_{E_{i}}|=\phi^{\epsilon,E}_{0},$ (3.9)
where $\phi^{\epsilon,E}_{0}$ admits for $\omega_{\epsilon}\mathrel{\mathop{\ordinarycolon}}=\omega-\epsilon\cdot 2\pi\cdot\sum_{i=1}^{r}a_{i}\cdot c_{1}(\mathscr{O}_{X}(E_{i}),h^{E_{i}})$ the following description
$\phi^{\epsilon,E}_{0}\mathrel{\mathop{\ordinarycolon}}=\sup\Big{\{}\phi\in{\textrm{PSH}}(X,\omega_{\epsilon})\mathrel{\mathop{\ordinarycolon}}\phi\leq-\epsilon\cdot\sum_{i=1}^{r}a_{i}\cdot\log|s_{E_{i}}|\Big{\}}.$ (3.10)
Note that in [2, Lemma 4.1], it was assumed that $E$ is irreducible, but the proof holds without this assumption. Note that our assumption on $\epsilon>0$ implies that ${\textrm{PSH}}(X,\omega_{\epsilon})$ is non-empty.
Now, let $K$ be an arbitrary non-pluripolar subset not intersecting $E$ (for example, a small ball in some local coordinates around $x\notin E$). Then for $C\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in K}(-\epsilon\cdot\sum_{i=1}^{r}a_{i}\cdot\log|s_{E_{i}}(x)|)$, we have $C<+\infty$. Let us now define $\phi_{K}\mathrel{\mathop{\ordinarycolon}}=\sup\{\phi\in{\textrm{PSH}}(X,\omega_{\epsilon})\mathrel{\mathop{\ordinarycolon}}\phi\leq C\text{ on }K\}$. The non-pluripolarity of $K$ and non-emptiness of ${\textrm{PSH}}(X,\omega_{\epsilon})$ assure that $\phi_{K}$ is uniformly bounded from above, cf. [19, Theorem 9.17]. The bound of Lemma 3.3 follows immediately from this, (3.9) and the obvious bound $\phi^{\epsilon,E}_{0}\leq\phi_{K}$. ∎
With all these preliminaries, we can finally complete our proof.
Proof of Theorem 2.3..
Note first that it suffices to establish the result for $\epsilon>0$ small enough. From now on, we choose $\epsilon$ as in Lemma 3.3. Remark first the following characterization of the partial Christoffel-Darboux kernel
$B_{k}^{\epsilon,Y}(x,x)=\sup\frac{|s(x)|_{(h^{L})^{k}}}{\|s\|_{{\textrm{Hilb}}_{k}(h^{L},\mu)}},$ (3.11)
where the supremum is taken over all $s\in H^{0}(X,L^{\otimes k}\otimes\mathscr{O}_{X}(-\lceil\epsilon kE\rceil))\setminus\{0\}$. Now, according to Lemma 3.1, there is a continuous metric $h^{L}_{0}$ on $L$, so that for any $k\in\mathbb{N}$, we have
$\|s\|_{{\textrm{Hilb}}_{k}(h^{L},\mu)}\geq\|s\|_{{\textrm{Ban}}_{k}^{\infty}(h^{L}_{0})}.$ (3.12)
By Lemma 3.2, for any $s\in H^{0}(X,L^{\otimes k}\otimes\mathscr{O}_{X}(-\lceil\epsilon kE\rceil))\setminus\{0\}$, we have
$\|s\|_{{\textrm{Ban}}_{k}^{\infty}(h^{L}_{0})}=\|s\|_{{\textrm{Ban}}_{k}^{\infty}(h^{L}_{0}\cdot\exp(-\phi^{\epsilon,E}_{0}))},$ (3.13)
where $\phi^{\epsilon,E}_{0}$ is defined analogously to (3.7) where we take a reference metric $h^{L}_{0}$ instead of $h^{L}$. By the definition of the $L^{\infty}$-norm, we have
$\|s\|_{{\textrm{Ban}}_{k}^{\infty}(X,h^{L}_{0}\cdot\exp(-\phi^{\epsilon,E}_{0}))}\geq|s(x)|_{(h^{L}_{0})^{k}}\cdot\exp(-k\cdot\phi^{\epsilon,E}_{0}(x))).$ (3.14)
Hence, from (3.11)-(3.14), we deduce that there is $C>0$, so that for any $x\in X$, we have
$B_{k}^{\epsilon,Y}(x,x)\leq\exp(Ck+k\cdot\phi^{\epsilon,E}_{0}(x)).$ (3.15)
The result then follows immediately from this and Lemma 3.3. ∎
4 Reduction to measures supported away from subvarieties
The main goal of this section is to show that if a given measure is truncated around a subvariety, then the associated Christoffel-Darboux kernel does not change much. As a consequence of this, we establish that in order to show Theorem 1.1 in full generality, it suffices to establish it solely for the measures supported away from a fixed subvariety.
It turns out that this form of “continuity under truncation” extends beyond subvarieties to a broader class – namely, those subsets that are not charged by any subsequential limit of the diagonal Christoffel-Darboux measures. For potential future applications, we therefore state our results in this more general setting.
More precisely, we fix a closed subset $A\subset X$ and consider a decreasing sequence of open neighborhoods $U_{1}\Supset U_{2}\Supset\cdots$ of $A$ such that $\bigcap_{i=1}^{+\infty}U_{i}=A$. Such a sequence always exists since $A$ is closed. By Urysohn’s lemma, we consider bump functions $\rho_{i}\mathrel{\mathop{\ordinarycolon}}X\to[0,1]$, $i\in\mathbb{N}^{*}$, verifying $\rho_{i}=1$ over $X\setminus U_{i}$ and $\rho_{i}=0$ over $U_{i+1}$. We assume that for any $\epsilon>0$, there is $i\in\mathbb{N}^{*}$, so that for any $k\in\mathbb{N}$, we have
$\int_{U_{i}}\mu_{k}^{\rm{Berg},\Delta}\leq\epsilon.$ (4.1)
We denote $\mu_{i}\mathrel{\mathop{\ordinarycolon}}=\rho_{i}\cdot\mu$, and let $B_{k,i}(x,y)\in L^{k}_{x}\otimes(L^{k}_{y})^{*}$, be the Christoffel-Darboux kernel associated with $h^{L}$ and $\mu_{i}$. We denote by $\mu_{k,i}^{\rm{Berg}}$ the associated probability measures on $X\times X$, defined as in (1.5). The main result of this section goes as follows.
Proposition 4.1.
Assume that a compact subset $K\subset X\times X$ is such that for any $\epsilon>0$, $i\in\mathbb{N}$, there is $k_{i}\in\mathbb{N}$, so that $\int_{K}\mu_{k,i}^{\rm{Berg}}\leq\epsilon$, for any $k\geq k_{i}$. Then for any $\epsilon>0$, there is $k_{0}\in\mathbb{N}$, so that $\int_{K}\mu_{k}^{\rm{Berg}}\leq\epsilon$, for any $k\geq k_{0}$.
In order to establish Proposition 4.1, the following lemma will be crucial.
Lemma 4.2.
For any $\epsilon>0$, there are $k_{0},i_{0}\in\mathbb{N}$, so that for any $k\geq k_{0}$ and $i\geq i_{0}$, we have
$\iint\Big{|}B_{k,i}(x,y)-B_{k}(x,y)\Big{|}_{(h^{L})^{k}}^{2}\cdot d\mu_{i}(x)d\mu_{i}(y)\leq\epsilon\cdot n_{k}.$ (4.2)
The proof of Lemma 4.2 is based on the following comparison result on the off-diagonal Christoffel-Darboux kernels due to Lubinsky [22, (3.3)].
Lemma 4.3.
Assume that the measures $\mu_{1}$ and $\mu_{2}$ are ordered as $\mu_{1}\leq\mu_{2}$. Then for any $x\in X$, the associated Christoffel-Darboux kernels are related as
$\int\Big{|}B_{k,1}(x,y)-B_{k,2}(x,y)\Big{|}_{(h^{L})^{k}}^{2}\cdot d\mu_{1}(y)\leq B_{k,1}(x,x)-B_{k,2}(x,x).$ (4.3)
Remark 4.4.
a) Observe that the characterization of the Christoffel-Darboux kernel analogous to (3.11) directly ensures that the right-hand side of (4.3) is positive.
b) Although Lubinsky in [22] establishes the above bound only in the setting of measures on the real line, the argument relies exclusively on the reproducing property of the Christoffel-Darboux kernel, and so it remains valid in our generality.
Proof of Lemma 4.2.
Note first that immediately from our construction of $\rho_{i}$, we have $\mu_{1}\leq\mu_{2}\leq\cdots\leq\mu$. From this and Lemma 4.3, we deduce that for any $i\in\mathbb{N}$, we have
$\iint\Big{|}B_{k,i}(x,y)-B_{k}(x,y)\Big{|}_{(h^{L})^{k}}^{2}\cdot d\mu_{i}(x)d\mu_{i}(y)\leq\int\Big{(}B_{k,i}(x,x)-B_{k}(x,x)\Big{)}\cdot d\mu_{i}(x).$ (4.4)
By our definition of $\mu_{i}$, we have the following bound
$\int B_{k}(x,x)\cdot d\mu_{i}(x)\geq\int_{X\setminus U_{i}}B_{k}(x,x)\cdot d\mu(x).$ (4.5)
By (2.7) and (4.1), for any $\epsilon>0$, there are $i_{0},k_{0}\in\mathbb{N}$, so that for any $i\geq i_{0}$, $k\geq k_{0}$,
$\int_{X\setminus U_{i}}B_{k}(x,x)\cdot d\mu(x)\geq(1-\epsilon)n_{k}.$ (4.6)
Moreover, it is immediate that
$\int B_{k,i}(x,x)\cdot d\mu_{i}(x)=n_{k}.$ (4.7)
Lemma 4.2 now follows immediately from (4.4), (4.5), (4.6) and (4.7). ∎
Proof of Proposition 4.1.
Directly from (4.1), we see that for any $\epsilon>0$, there are $i_{0},k_{0}\in\mathbb{N}$, so that for any $i\geq i_{0}$, $k\geq k_{0}$, we have
$\iint_{U_{i}\times X}\big{|}B_{k}(x,y)\big{|}_{(h^{L})^{k}}^{2}\cdot d\mu(x)d\mu(y)\leq\frac{\epsilon\cdot n_{k}}{4}.$ (4.8)
Moreover, from Lemma 4.2, our assumption $\int_{K}\mu_{k,i}^{\rm{Berg}}\leq\epsilon$, and Minkowski inequality, we see that for any $\epsilon>0$, there are $k_{0}\in\mathbb{N}$, $i\geq i_{0}$, so that for any $k\geq k_{0}$, we have
$\iint_{K}\big{|}B_{k}(x,y)\big{|}_{(h^{L})^{k}}^{2}\cdot d\mu_{i}(x)d\mu_{i}(y)\leq\frac{\epsilon\cdot n_{k}}{2}.$ (4.9)
Note, however, that by the definition of $\mu_{i}$, we have
$\iint_{K}\big{|}B_{k}(x,y)\big{|}_{(h^{L})^{k}}^{2}\cdot d\mu_{i}(x)d\mu_{i}(y)\geq\iint_{K}\big{|}B_{k}(x,y)\big{|}_{(h^{L})^{k}}^{2}\cdot d\mu(x)d\mu(y)\\ -2\cdot\iint_{U_{i}\times X}\big{|}B_{k}(x,y)\big{|}_{(h^{L})^{k}}^{2}\cdot d\mu(x)d\mu(y).$ (4.10)
A combination of (4.8), (4.9) and (4.10) yields Proposition 4.1. ∎
5 Decay of the kernel for measures supported away from a divisor
The main goal of this section section is to show that for measures supported away from some divisors, a quantitative version of Theorem 1.1 can be established. As a consequence of this and the preceding analysis, we then deduce Theorem 1.1 in full generality.
We will now assume that $k_{0}\in\mathbb{N}$ is big enough so that $n_{k_{0}}\geq 2$. We fix two linearly independent sections $s_{1},s_{2}\in H^{0}(X,L^{\otimes k_{0}})$, and consider the subvariety $Y\mathrel{\mathop{\ordinarycolon}}={\textrm{div}}(s_{2})$. We fix a Borel measure $\mu_{0}$ on $X$ whose support is disjoint from $Y$. Assume, moreover, that the support of $\mu_{0}$ is not an analytic subset. Then the form defined as in (1.1), but for $\mu_{0}$ instead of $\mu$, is positive definite, and hence the associated Christoffel-Darboux kernel, $B_{k,0}(x,y)\in L^{k}_{x}\otimes(L^{k}_{y})^{*}$, can be defined. The main result of this section goes as follows.
Theorem 5.1.
There is $C>0$, so that for any $k\in\mathbb{N}^{*}$, we have
$\iint\big{|}B_{k,0}(x,y)\big{|}^{2}_{(h^{L})^{k}}\cdot\Big{|}\frac{s_{1}(x)}{s_{2}(x)}-\frac{s_{1}(y)}{s_{2}(y)}\Big{|}^{2}\cdot d\mu_{0}(x)d\mu_{0}(y)\leq C\cdot(n_{k}-n_{k-k_{0}}).$ (5.1)
Remark 5.2.
a) Note that by Fujita’s theorem [16], recalled after (2.3), the right-hand side of (5.1) is asymptotically smaller than $n_{k}$. In particular, if for a compact subset $K\subset(X\setminus Y)\times(X\setminus Y)$, there exists a constant $C_{1}>0$ such that $|s_{1}(x)/s_{2}(x)-s_{1}(y)/s_{2}(y)|\geq C_{1}$, for any $x\times y\in K$, then Theorem 5.1 provides a quantitative version of Theorem 1.1.
b) In the setting of orthogonal polynomials described in (1.4), one can consider $s_{2}\mathrel{\mathop{\ordinarycolon}}=\sigma^{2}$ and $s_{1}=z\cdot\sigma^{2}$, where $\sigma\in H^{0}(\mathbb{P}^{1},\mathscr{O}(1))$ is as described after (1.4). Then for any compact $K\subset\mathbb{C}\times\mathbb{C}\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$, not intersecting the diagonal, the above $C_{1}$ exists. We conclude that for arbitrary compactly supported measures $\mu$ on $\mathbb{C}$ with non-finite support, the estimate (1.7) holds. We note that when the support of $\mu$ lies inside of the real axis, one can give a shorter proof of (1.7) based on the Christoffel-Darboux formula, see in particular [9, (2.13)].
Proof.
Our proof is inspired by the the proof of [4, Theorem 17] by Berman-Ortega-Cerdá in the sense that the estimate (5.1) is established by studying the Hilbert-Schmidt norm of a certain sequence of operators. However, in contrast to [4], our operators are not of Toeplitz type. This allows us to bypass – at the expense of a more delicate analysis of the associated Schwartz kernel – the assumption present in [4] that the support of the measure is totally real.
More precisely, we let $f\mathrel{\mathop{\ordinarycolon}}=s_{1}/s_{2}$ and consider a sequence of operators $S_{k}\mathrel{\mathop{\ordinarycolon}}L^{2}(X,L^{\otimes k},\mu_{0})\to L^{2}(X,L^{\otimes k},\mu_{0})$, defined through their Schwartz kernels
$S_{k}(x,y)\mathrel{\mathop{\ordinarycolon}}=\int B_{k,0}(x,z)\cdot\big{(}f(x)-f(z)\big{)}\cdot B_{k,0}(z,y)\cdot d\mu_{0}(z).$ (5.2)
Let us verify that the operators $S_{k}$, $k\in\mathbb{N}$, are uniformly bounded, i.e. there is $C>0$, so that $\|S_{k}\|\leq C$, where $\|\cdot\|$ is the operator norm. Since the measure $\mu_{0}$ is supported away from $Y$, the function $f$ is uniformly bounded on its support. The operator $S_{k}$ is given by a composition of a Bergman projection (which is of unit norm) and the commutator of the Bergman projection and the multiplication operator by $f$ (which is – by the argument above – of bounded norm, independent of $k\in\mathbb{N}$). Hence, there is $C>0$, so that $\|S_{k}\|\leq C$, for any $k\in\mathbb{N}$.
Now, let us verify that the kernel of $S_{k}$ has codimension at most $n_{k}-n_{k-k_{0}}$. Since the second factor of $S_{k}$ is given by the Bergman projection, we see that $H^{0}(X,L^{\otimes k})^{\perp}\subset\ker S_{k}$. Now, consider a sequence of linear operators $F_{k}\mathrel{\mathop{\ordinarycolon}}H^{0}(X,L^{\otimes(k-k_{0})})\to H^{0}(X,L^{\otimes k})$, given by a multiplication by $s_{2}$. It is immediate to verify that $\operatorname{Im}F_{k}\subset\ker S_{k}$. But the map $F_{k}$ is clearly an injection, and so the dimension of its image is $n_{k-k_{0}}$. This along with the obvious fact $\operatorname{Im}F_{k}\cap H^{0}(X,L^{\otimes k})^{\perp}=\emptyset$ yields the needed bound on the codimension of the kernel of $S_{k}$.
As a conclusion, we deduce the following bound
$\|S_{k}\|_{{\textrm{HS}}}^{2}\leq C\cdot(n_{k}-n_{k-k_{0}}),$ (5.3)
where $\|\cdot\|_{{\textrm{HS}}}$ is the Hilbert-Schmidt norm. Since the Hilbert-Schmidt norm of an integral operator can be expressed through the $L^{2}$-norm of its Schwartz kernel, we deduce
$\iint\big{|}S_{k}(x,y)\big{|}^{2}_{(h^{L})^{k}}\cdot d\mu_{0}(x)d\mu_{0}(y)\leq C\cdot(n_{k}-n_{k-k_{0}}).$ (5.4)
We shall now see that (5.4) is essentially a restatement of (5.1). By using the identity $\big{|}S_{k}(x,y)\big{|}^{2}_{(h^{L})^{k}}=S_{k}(x,y)\cdot S_{k}(x,y)^{*}$ and (5.2), we rewrite
$\iint\big{|}S_{k}(x,y)\big{|}^{2}_{(h^{L})^{k}}\cdot d\mu_{0}(x)d\mu_{0}(y)\\ =\iint\int B_{k,0}(x,z)\cdot\big{(}f(x)-f(z)\big{)}\cdot B_{k,0}(z,y)\cdot d\mu_{0}(z)\cdot\\ \cdot\int B_{k,0}(x,w)^{*}\cdot\big{(}\overline{f}(x)-\overline{f}(w)\big{)}\cdot B_{k,0}(w,y)^{*}\cdot d\mu_{0}(w)\cdot d\mu_{0}(x)d\mu_{0}(y).$ (5.5)
Using Fubini’s theorem and the symmetry $B_{k,0}(x,y)^{*}=B_{k,0}(y,x)$, we further rewrite
$\iint\big{|}S_{k}(x,y)\big{|}^{2}_{(h^{L})^{k}}\cdot d\mu_{0}(x)d\mu_{0}(y)\\ =\iiiint B_{k,0}(x,z)B_{k,0}(z,y)B_{k,0}(w,x)B_{k,0}(y,w)\cdot\\ \cdot\big{(}f(x)-f(z)\big{)}\cdot\big{(}\overline{f}(x)-\overline{f}(w)\big{)}\cdot d\mu_{0}(z)d\mu_{0}(w)d\mu_{0}(x)d\mu_{0}(y).$ (5.6)
We expand the integrand, apply the reproducing property of the Bergman kernel to eliminate two of the four kernel factors in each of the resulting terms, and finally rename the coordinates to obtain:
$\iint\big{|}S_{k}(x,y)\big{|}^{2}_{(h^{L})^{k}}\cdot d\mu_{0}(x)d\mu_{0}(y)=\frac{1}{2}\iint\big{|}B_{k,0}(x,y)\big{|}^{2}_{(h^{L})^{k}}\cdot\big{|}f(x)-f(y)\big{|}^{2}\cdot d\mu_{0}(x)d\mu_{0}(y),$ (5.7)
which finishes the proof of Theorem 5.1 by (5.4). ∎
We are now finally in a position to prove the main result of this paper.
Proof of Theorem 1.1.
[…]
We now consider a subset $Y\mathrel{\mathop{\ordinarycolon}}=\Sigma\cup{\textrm{div}}(s_{1})\cup\cdots\cup{\textrm{div}}(s_{n_{k}})$, where $s_{i}$, $i=1,\ldots,n_{k}$, is a fixed basis of $H^{0}(X,L^{\otimes k})$. The set $Y$ is subvariety. Hence, according to Theorem 2.1 and Proposition 4.1, it suffices to establish Theorem 1.1 for measures $\mu$ which are supported away from $Y$. Note, however, that for such measures $\mu$, for a fixed $K\subset X\times X$ not intersecting the diagonal, there is $C>0$, so that for $K^{\prime}\mathrel{\mathop{\ordinarycolon}}=K\cap{\rm{supp}}(\mu)\times{\rm{supp}}(\mu)$, for any $x,y\in K^{\prime}$, we have
$\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n_{k}}\Big{|}\frac{s_{i}(x)}{s_{j}(x)}-\frac{s_{i}(y)}{s_{j}(y)}\Big{|}^{2}\geq C.$ (5.8)
Theorem 1.1 for such measures $\mu$ now follows immediately from Theorem 5.1, Remark 5.2 and (5.8). As explained above, this finishes our proof. ∎
6 Pointwise localization of Christoffel-Darboux kernel
This section aims to present a problem closely related in spirit to the main result, but situated in the context of pointwise convergence. For a given $x\in X$, we introduce the following measure
$\mu_{k}^{x}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{B_{k}(x,x)}|B_{k}(x,y)|^{2}_{(h^{L})^{k}}\cdot d\mu(y).$ (6.1)
Immediately from (1.3), we see that $\mu_{k}^{x}$ is a probability measure. Due to the same heuristics as described after (1.3), one should expect that the measures $\mu_{k}^{x}$ converge weakly towards the Dirac mass at $x$, as $k\to\infty$, when $x\in{\textrm{supp}}(\mu)$. This naive expectation turns out to be false; for counterexamples, we refer the reader to Breuer-Last-Simon [9, §3-5]. Following [9], measures that nevertheless satisfy this condition are said to satisfy the Nevai condition at $x$.
It is natural to conjecture – especially in light of Theorem 1.1 – that the following statement, which extends [9, Conjecture 1.4] from compactly supported measures on $\mathbb{R}\subset\mathbb{P}^{1}$ to more general measures, might hold: Nevai condition holds for $\mu$-almost every $x$.
Note that this conjecture holds (in a much stronger form, for every point $x$) when $\mu$ is a volume form and $h^{L}$ is a strictly positive smooth metric, by the results of Dai-Liu-Ma [12], Christ [10], and Ma-Marinescu [25]. In the setting of orthogonal polynomials associated with a broad class of compactly supported measures on $\mathbb{R}$, the conjecture was established by Nevai [26] (again in the stronger sense, for every $x\in{\textrm{supp}}(\mu)$); see also [9].
Let us describe the connection between Theorem 1.1 and the above conjecture. For this, we introduce the following measures on $X$
$\nu_{k}^{x}\mathrel{\mathop{\ordinarycolon}}=\frac{B_{k}(x,x)}{n_{k}}\cdot\mu_{k}^{x}.$ (6.2)
Unlike $\mu_{k}^{x}$, the measures $\nu_{k}^{x}$ are not necessarily of unit mass. However, by (1.3), the whole ensemble of measures $\nu_{k}^{x}$, $x\in X$, is of unit mass on average, i.e.
$\int_{x\in X}{\rm{vol}}(\nu_{k}^{x})d\mu(x)=1,\quad\text{where}\quad{\rm{vol}}(\nu_{k}^{x})\mathrel{\mathop{\ordinarycolon}}=\int_{X}d\nu_{k}^{x}(x).$ (6.3)
Proposition 6.1.
For any increasing sequence $k_{i}\in\mathbb{N}$, $i\in\mathbb{N}$, there is an increasing subsequence $l_{i}$, $i\in\mathbb{N}$, so that for $\mu$-almost any $x\in X$, the following holds. For any $f\in\mathscr{C}^{0}(X)$, vanishing in a neighborhood of $x$, we have $\int_{X}f(y)\cdot d\nu_{l_{i}}^{x}(y)\to 0$, as $i\to\infty$.
Proof.
Theorem 1.1 states that for any compact subset $K\subset X\times X$ disjoint from the diagonal $\Delta$, the sequence of functions $g_{k}(x,y)\mathrel{\mathop{\ordinarycolon}}=\mathbf{1}_{K}(x,y)\cdot\frac{1}{n_{k}}\left|B_{k}(x,y)\right|^{2}_{(h^{L})^{k}}$ converges to zero in $L^{1}(X\times X,\mu\times\mu)$, as $k\to\infty$. It follows that, given any increasing sequence $k_{i}\in\mathbb{N}$, $i\in\mathbb{N}$, there exists a subsequence $l_{i}\in\mathbb{N}$, $i\in\mathbb{N}$, such that for $\mu$-almost every $x$, the functions $y\mapsto g_{l_{i}}(x,y)$ converge to 0 in $L^{1}(\mu)$, as $i\to\infty$. This establishes the desired result for functions $f$ verifying $(x,{\textrm{supp}}(f))\subset K$. To extend this to the general case, consider a nested sequence of compact subsets $K_{1}\subset K_{2}\subset\cdots\subset X\times X\setminus\Delta$ such that $\cup_{i=1}^{\infty}K_{i}=X\times X\setminus\Delta$. Applying the above argument to each $K_{i}$, we extract a subsequence $l^{i}_{j}\in\mathbb{N}$, $j\in\mathbb{N}$, for each $i$. An easy verification shows that the diagonal subsequence $l^{i}_{i}$, $i\in\mathbb{N}$, satisfies our claim. ∎
7 Applications to the theory of Toeplitz operators
The main goal of this section is to establish that spectral equidistribution of Toeplitz operators is equivalent to the convergence of (diagonal) Christoffel-Darboux measure. More precisely, the main result of this section goes as follows.
Theorem 7.1.
For any increasing sequence $k_{i}\in\mathbb{N}$, $i\in\mathbb{N}$, the following are equivalent:
  • The sequence $\mu_{k_{i}}^{\rm{Berg},\Delta}$ converges weakly, as $i\to\infty$, towards a measure $\nu$.
  • The sequence $\mu_{k_{i}}^{\rm{Berg}}$ converges weakly, as $i\to\infty$, towards the measure $\Delta_{*}\nu$, where $\Delta\mathrel{\mathop{\ordinarycolon}}X\to X\times X$ is the diagonal embedding.
  • For any continuous function $f\mathrel{\mathop{\ordinarycolon}}X\to\real$, the spectral measures $\mu_{k_{i}}^{Spec}(f)\mathrel{\mathop{\ordinarycolon}}=\frac{1}{n_{k_{i}}}\sum_{\lambda\in{\rm{Spec}}(T_{k_{i}}(f))}\delta[\lambda]$ of $T_{k_{i}}(f)$ converge weakly, as $i\to\infty$, towards a measure $f_{*}\nu$.
Proof.
An immediate verification shows that for any $f\mathrel{\mathop{\ordinarycolon}}X\to\real$, we have
$\int f(x)\cdot d\mu_{k_{i}}^{\rm{Berg},\Delta}(x)=\frac{1}{n_{k_{i}}}{\rm{Tr}}\big{[}T_{k_{i}}(f)\big{]}=\int x\cdot d\mu_{k_{i},f}^{Spec}(x).$ (7.1)
The implication $c)\Rightarrow a)$ is then immediate. Note also that for any of the two natural projections $\pi\mathrel{\mathop{\ordinarycolon}}X\times X\to X$, we easily deduce $\pi_{*}\mu_{k_{i}}^{\rm{Berg}}=\mu_{k_{i}}^{\rm{Berg},\Delta}$, and so $b)\Rightarrow a)$ is also immediate.
The proof of the reverse implications $a)\Rightarrow b)$, $a)\Rightarrow c)$ follows the same argument as in [14, Theorems 1.1 and 1.7], with a few adjustments that we detail below. First, note that [14] only treated the case of Bernstein-Markov measures and considered limits along the full sequence indexed by $\mathbb{N}$, rather than along subsequences. The passage from limits along $\mathbb{N}$ to limits along subsequences is immediate. Moreover, in the proof of [14, Theorems 1.1 and 1.7], the Bernstein-Markov assumption was only used to establish the validity of Corollary 1.3 (which was stated in [14] with that additional assumption) and to invoke the equidistribution result of Berman-Boucksom-Witt Nyström [3] concerning the diagonal Christoffel-Darboux measures $\mu_{k_{i}}^{\mathrm{Berg},\Delta}$. In the present article, we have shown that Corollary 1.3 remains valid without assuming the Bernstein-Markov property. Furthermore, in statement $a)$, we assume the equidistribution of the diagonal Christoffel-Darboux measures $\mu_{k_{i}}^{\mathrm{Berg},\Delta}$. It follows that the proof carries over without modification. ∎